Which of the following is not equal to the others? (A) 6/8 (B) 12/16 (C) 15/25 (D) 18/24

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

All the options given in the question are further simplified as,
(A) 6/8
Divide both numerator and denominator by 2.
= 3/4

Option (B) 12/16
Divide both numerator and denominator by 4. = 3/4
Option(C) 15/25
Divide both numerator and denominator by 5. = 3/5
Option(D) 18/24
Divide both numerator and denominator by 6.
=3⁄4
on Comparing all results,
(3⁄4 = 3⁄4 = 3⁄4)≠3/5
Thus, (6/8 = 12/16 = 18/24)≠15/25



Thanks